OpenStudy (anonymous):

[8.03] What are the factors of x2 − 3x − 28? (x + 7)(x − 4) (x − 7)(x + 4) (x + 2)(x − 14) (x − 2)(x + 14)

OpenStudy (anonymous):



OpenStudy (anonymous):

Wait, isn't it b?

OpenStudy (anonymous):

yeah ,sorry it s b

OpenStudy (anonymous):

(x-7)(x+4)

OpenStudy (anonymous):

Oh, ok, thx, lol, anbother one?

OpenStudy (anonymous):

yep lol

OpenStudy (anonymous):

8.07] Jose factored 27x2 + 48y2 using the following steps. 27x2 + 48y2 = 3(9x2 + 16y2) = 3(3x + 4y)(3x − 4y) Determine whether Jose factored this binomial correctly. If not, use complete sentences to explain what he did wrong and how he could check his factors for accuracy in the future.

OpenStudy (anonymous):

BTW, Thx, lol! I need this. :)

OpenStudy (anonymous):

it is wrong

OpenStudy (anonymous):

Oh, okay, thx, how?

OpenStudy (anonymous):

I agree, with his calc it would return a value of 27x2 - 48y2

OpenStudy (anonymous):

Can you tell me how?

OpenStudy (anonymous):

(3x+4y)(3x-4y) 9x2 - 12xy + 12xy -16y2 Then simplify 9x2 - 16y2

OpenStudy (anonymous):

OMG, thx so much, you guys are so helpful!!!!! :))) ♥♥♥

OpenStudy (anonymous):